OpenStudy (manman):

How to Solve this by the Elimination method Step by Step 2x-3y=8 3x-2y=15

OpenStudy (ajprincess):

2x-3y=18-(1) 3x-2y=15-(2) Multiply (1) by 2 nd multiply (2) by 3. Then subtract the eqn u get by multiplying (2) by 3 from the eqn u get by multiplying (1) by 3. U will get an equation only in terms of x. Solve for x. Then substituting the value of x in either (1) or (2) find y.